Two mutually perpendicular oscillations
have angular frequency o)0 and amplitudes
al and a2 such that al > a2. If the initial
phase difference between these is 4),
determine the nature of the resultant
oscillation obtained on their superposition.
1
Expert's answer
2019-04-30T09:53:27-0400

The equation of the resultant path is given by formula


"\frac {x^2} {a_1 ^2} + \frac {y^2} {a_2 ^2} – 2 \frac {xy} {a_1 a_2} \cos \phi = \sin^2 \phi (1)"

In our case,

"\phi=4\pi"

Using (1) we got:


"\frac {x^2} {a_1 ^2} + \frac {y^2} {a_2 ^2} – 2 \frac {xy} {a_1 a_2} = 0"

Answer:


"\frac {x^2} {a_1 ^2} + \frac {y^2} {a_2 ^2} – 2 \frac {xy} {a_1 a_2} = 0"